OpenStudy (anonymous):

2x - y = 5 x + 3y = 7 What is the second row of entries for the system determinant? 0, 3 1, 3 3, 7

OpenStudy (anonymous):

These are the options? :/ I mean we have to put them in the equation instead of x and y?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

What? :/ @wtbryant i was actually asking you :/ I didn't answer the question :/

OpenStudy (anonymous):

oh sorry yeah those are the options

OpenStudy (anonymous):

Um okay. But all the options are wrong :/

OpenStudy (anonymous):

It will be none of the above

OpenStudy (anonymous):

yea i figured tht out im stuck

OpenStudy (anonymous):

what

OpenStudy (anonymous):

I mean such an easy question is this why you are struck :/

OpenStudy (anonymous):

stuck*

OpenStudy (anonymous):

yea i guess

OpenStudy (anonymous):

You have only these 3 options?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i cant figure this out

OpenStudy (anonymous):

None of the above is the answer :)

OpenStudy (anonymous):

but that is not a coice

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

2x - y = 5 x + 3y = 7 now put x=0 and y=3 Solve it and show :)

OpenStudy (anonymous):

2(0)-(3)=-3

OpenStudy (anonymous):

(0)+3(3)=9

OpenStudy (anonymous):

2x - y = 5 x + 3y = 7 2(0)-(3)=5 0-3=-3 0+3(3)=7 0+9=7 9=7 So this is wrong because left hand side is not equal to right hand side :)

OpenStudy (anonymous):

YEA

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Solve like this how I did and show :) Now put x=1, y= 3

OpenStudy (anonymous):

(1)+3(3)=5 1+9=10 5=10

OpenStudy (anonymous):

Now put x=3, y= 7

OpenStudy (anonymous):

2(1)-(3)=7 2-3=-1 7=-1

OpenStudy (anonymous):

(3)+3(7)=7 3+21=24 7=24

OpenStudy (anonymous):

So all these options are wrong :/

OpenStudy (anonymous):

yea i guess

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks for your help though

OpenStudy (anonymous):

Your welcome :)